# Climbing the Ladder of Interpretability with Counterfactual Concept Bottleneck Models

### v1

@@ Abstract
Current deep learning models are not designed to simultaneously address three fundamental questions: predict class labels to solve a given classification task (the “What?”), explain task predictions (the “Why?”), and imagine alternative scenarios that could result in different predictions (the “What if?”). The inability to answer these questions represents a crucial gap in deploying reliable AI agents, calibrating human trust, and deepening human-machine interaction. To bridge this gap, we introduce CounterFactual Concept Bottleneck Models (CF-CBMs), a class of models designed to efficiently address the above queries all at once without the need to run post-hoc searches. Our results show that CF-CBMs produce: accurate predictions (the “What?”), simple explanations for task predictions (the “Why?”), and interpretable counterfactuals (the “What if?”). CF-CBMs can also sample or estimate the most probable counterfactual to: (i) explain the effect of concept interventions on tasks, (ii) show users how to get a desired class label, and (iii) propose concept interventions via “task-driven” interventions.
@@ 1 Introduction
To calibrate human trust and enhance human-machine interactions, deep learning (DL) models should learn how to master three fundamental questions: predict class labels for new inputs (the “What?”), explain predictions (the “Why?”), and imagine alternative scenarios that could result in different predictions (the “What if?”)—also known as counterfactual explanations (Wachter et al., 2017). We refer to this escalating sequence of questions as the “ladder of interpretability” (Figure 1), taking inspiration from a similar analogy proposed by Pearl et al. (2016). Crucially, no DL model is currently designed to cover the whole ladder.
Figure 1: CF-CBMs are designed to address three fundamental questions represented by the “ladder of interpretability”: 1. predict class labels for a given task (“What?”), 2. explain predictions via learned concepts (“Why?”), and 3. imagine concept-based counterfactuals that would have led to different class predictions (“What if?”).
At the base of the ladder, standard DL models excel at addressing “what?” questions. Standing on the second rung, Concept Bottleneck Models (CBMs) (Koh et al., 2020) currently represent a higher-ranked DL model in this ladder, as they are designed to answer both “what?” and “why?” questions. To this end, CBMs learn to explain their own predictions by first predicting a set of human-understandable concepts (Kim et al., 2018) and then using the learned concepts to solve the task. Yet, their design and training do not extend to “what if?” queries, leaving the ladder’s third rung still unreachable for these models.
Contributions.
To bridge this gap, we introduce CounterFactual Concept Bottleneck Models (CF-CBM), the first CBMs designed to cover the entire ladder of interpretability. Our key innovation lies in a latent process generating two similar concept vectors. The model uses the first concept vector to predict the target class label (similar to standard CBMs), while the second one is used to predict a different class label. Our experimental results show that CF-CBMs generate: (“What?”) accurate concept/class predictions, (“Why?”) simpler and more robust explanations compared to standard CBMs, (“What if?”) interpretable concept-based counterfactuals. Compared to standard CBMs, CF-CBMs can also sample or estimate the most probable counterfactual to: (i) explain the effect of concept interventions on the classification task, (ii) show users how to get a desired class label, and (iii) propose concept interventions via “task-driven” interventions. The code is available at https://github.com/gabriele-dominici/Counterfactual-CBM.
@@
Figure 2: Counterfactual CBMs inference. For a given input sample, the task predictor answers “what?” queries predicting class labels. The concept predictor answers “why?” queries explaining class predictions via concept labels when taking as input the latent concept distribution $z$. When taking the counterfactual latent distribution $z^{\prime}$, the concept predictor answers “what if?” queries via concept-based counterfactuals.
@@ 2 Background
Concept Bottleneck Models (CBMs).
A CBM (Koh et al., 2020; Poeta et al., 2023) is a function composed by: (i) a concept encoder $g:X\to C$ mapping each entity $x\in X\subseteq\mathbb{R}^{d}$ (e.g., an image) to a set of $k$ concepts $\hat{c}\in C\subseteq[0,1]^{k}$ (e.g., “red”, “round”), and (ii) an interpretable (Rudin, 2019) task predictor $f:C\to Y$ mapping concepts to a set of $m$ classes $\hat{y}\in Y\subseteq[0,1]^{m}$ (e.g., “apple”) representing a downstream task. Each component $g_{i}$ and $f_{j}$ denotes the prediction score of the $i$-th concept and $j$-th class, respectively. During training, concept and class predictions $(\hat{c},\hat{y})$ are aligned with ground-truth labels $(c,y)$. At test time, a CBM provides explanations for class predictions indicating the presence or absence of a concept, and human experts may intervene on mispredicted concept labels to improve CBMs’ task performance. In a probabilistic perspective (Bahadori and Heckerman, 2020; Misino et al., 2022), (ideal) concepts $c$ represent key factors of variation (Kingma and Welling, 2013; Zarlenga et al., 2023) of observed data $x$ and $y$, as illustrated in the graphical model $x$$c$$y$
Counterfactual Explanations.
A counterfactual is a hypothetical statement in contrast with actual events that helps us understand the potential consequences of different choices or circumstances  (Pearl et al., 2016). In the context of machine learning, Wachter et al. (2017) define counterfactual explanations as an optimization problem representing an answer to a “what if” question. The optimization aims to find for each observation $x$ the closest datapoint $x^{\prime}$ such that a classifier $m:X\rightarrow Y$ produces a class label $m(x^{\prime})$ that is different from the original label $m(x)$.
@@ 3 Counterfactual Concept Bottleneck Models
To cover the ladder of interpretability, we propose CounterFactual Concept Bottleneck Models (CF-CBMs, Figure 2), a novel class of interpretable models designed to generate counterfactuals via variational inference. In the following, we motivate CF-CBMs’ design (Sec. 3.1), architecture and optimization objective (Sec. 3.2). Sections 3.4 and 3.3 show how CF-CBMs climb the ladder of interpretability and how users can act upon generated counterfactuals.
3.1 Motivation
We aim to design a model that generates interpretable “concept-based” counterfactuals i.e., counterfactuals enabling humans to act upon meaningful and interpretable units of information (Abid et al., 2022). For instance, “change the intensity of pixel 42 from 0 to 1” does not provide a clear (actionable) explanation of how to transform a “banana” to an “apple” for a human. In contrast, “change the shape of the object to round” provides an interpretable abstraction in terms of human-understandable concepts (Kim et al., 2018). In general, whenever the input features are too fine-grained (e.g., pixel intensities), the corresponding counterfactuals might not follow human semantics (Kim et al., 2018). Concept-based counterfactuals, instead, provide meaningful insights for users to act upon. This makes CF-CBM an ideal architecture to cover the three rungs of the ladder of interpretability.
@@ 3.2 Generating Concept-Based Counterfactuals
We also notice that multiple counterfactuals can be used to explain any given fact (Wachter et al., 2017; Pearl, 2000). To address this, an ideal solution would be to use a generative approach modeling the discrete distributions over $c$ and $c^{\prime}$. However, modeling such distributions directly might be unfeasible in practice (Richardson and Domingos, 2006) as it requires either modelling complex concept dependencies with discrete distributions which scale exponentially with the number of concepts, or assuming that concepts are independent of each other, an often unrealistic assumption. For this reason, we use a latent variable approach (Kingma and Welling, 2013) to model concept dependencies in a continuous latent distribution, which allows us to model discrete concept values as independent of each other given the latent variable.
Architecture.
CF-CBMs are latent variable models generating counterfactuals via variational inference. To this end, they add two random variables $z$ and $z^{\prime}$ to Diagram 1. These variables represent latent factors of variation whose probability distributions are easier to model and sample compared to those for $c$ and $c^{\prime}$. We also include arrows from $c$ and $y$ to the counterfactual latent distribution $z^{\prime}$ in order to explicitly model the dependency of $z^{\prime}$ on symbolic values of actual concept and class labels, resulting in the following overall probabilistic graphical model:
@@
$p(c,y|z)=p(y|c)p(c|z)$ (4)
$p(c^{\prime},y^{\prime}|z^{\prime})=p(y^{\prime}|c^{\prime})p(c^{\prime}|z^{\prime})$ (5)
$p(\mathbf{z}|c,y)=p(z)p(z^{\prime}|z,c,y)$ (6)
In our approach, $p(y|c)$ and $p(y^{\prime}|c^{\prime})$ are modeled as categorical distributions parametrized by the same task predictor $f$; $p(c|z)$ and $p(c^{\prime}|z^{\prime})$ as sets of Bernoulli distributions parametrized by the same concept predictor $\phi_{z}$. In practice, we assume that the input $x$ is always observed at test time, making the term and $p(x|z)$ irrelevant. Finally, $p(z)$ is a standard normal prior distribution and $p(z^{\prime}|z,c,y)$ is a learnable normal prior whose mean and variance are parametrized by a pair of neural networks $\phi_{p\mu}$ and $\phi_{p\sigma}$.
Amortized inference.
@@
CF-CBMs amortize inference needed for training by introducing two approximate Gaussian posteriors $q(z|x)$ and $q(z^{\prime}|z,c,y,y^{\prime})$ whose mean and variance are parametrized by a pair of neural networks $(\phi_{\mu},\phi_{\sigma})$ ($(\phi_{\mu^{\prime}},\phi_{\sigma^{\prime}})$, respectively). The corresponding inference graphical model (i.e. the encoder) is shown in Appendix A.
Optimization problem.
@@
CF-CBMs are trained to maximize the log-likelihood of tuples $(c,y,y^{\prime})$, while observing $x$. Following a variational inference approach, we optimize the evidence lower bound of the log-likelihood, which results (see Appendix A) in the following objective function to maximize:
@@
$\mathcal{L}=$ $\overbrace{\mathbb{E}_{z\sim q(z|x)}[\log p(c|z)]+\log p(y|c)}^{\text{ reconstruction of $c$ and $y$}}-\overbrace{D_{KL}[q(z|x)||p(z)]}^{\text{ prior regularization on $z$}}$
@@
$+$ $\overbrace{\mathbb{E}_{z,z^{\prime},c^{\prime}\sim p(c^{\prime}|z^{\prime})q(z^{\prime}|\alpha)q(z|x))}[\log p(y^{\prime}|c^{\prime})]}^{\text{reconstruction of $y^{\prime}$}}$
$-\overbrace{D_{KL}[q(z^{\prime}|\alpha)||p(z^{\prime}|z,c,y)]}^{\text{prior regularization on $z^{\prime}$ }}$
@@
where $D_{KL}$ is the Kullback–Leibler divergence and $\alpha=(z,c,y,y^{\prime})$. Moreover, in order to enforce concept-based counterfactuals to be as close as possible to the current concept labels, we add an additional term to the objective:
3.3 Climbing the “Ladder of Interpretability”
@@
We claim that CF-CBMs climb the ladder of interpretability by design, answering “what?”, “why?” and “what if?” queries through the following steps (Figure 2):
  1. 1.
    What? predict concept and class labels:
@@
      Sample from latent concept posterior:
      $z\sim\mathcal{N}(\phi_{\mu}(x),\phi_{\sigma}(x))$
@@
      Sample to predict concept and class labels:
      $\hat{y}\sim\text{Categorical}(f(\hat{c})),\ \hat{c}\sim\text{Bernoulli}(\phi_{z}(z))$
  2. 2.
    Why? explain class predictions using the interpretable model $f$ which identifies relevant concepts.
  3. 3.
    What if? imagine interpretable counterfactuals via the conditional generation of new concept tuples:

### v3

# Counterfactual Concept Bottleneck Models

@@ Abstract
Current deep learning models are not designed to simultaneously address three fundamental questions: predict class labels to solve a given classification task (the "What?"), simulate changes in the situation to evaluate how this impacts class predictions (the "How?"), and imagine how the scenario should change to result in different class predictions (the "Why not?"). While current approaches in causal representation learning and concept interpretability are designed to address some of these questions individually (such as Concept Bottleneck Models, which address both “what” and “how” questions), no current deep learning model is specifically built to answer all of them at the same time. To bridge this gap, we introduce CounterFactual Concept Bottleneck Models (CF-CBMs), a class of models designed to efficiently address the above queries all at once without the need to run post-hoc searches. Our experimental results demonstrate that CF-CBMs: achieve classification accuracy comparable to black-box models and existing CBMs (“What?”), rely on fewer important concepts leading to simpler explanations (“How?”), and produce interpretable, concept-based counterfactuals (“Why not?”). Additionally, we show that training the counterfactual generator jointly with the CBM leads to two key improvements: (i) it alters the model’s decision-making process, making the model rely on fewer important concepts (leading to simpler explanations), and (ii) it significantly increases the causal effect of concept interventions on class predictions, making the model more responsive to these changes.
@@ 1 Introduction
Figure 1: CF-CBMs generate counterfactuals at the concept level rather than at the input level, as the changes are more interpretable. On the contrary, identifying the changes in the input level require significant more effort from the user. The images are sourced from the SIIM Pneumothorax dataset (Zawacki et al., 2019).
@@
To calibrate human trust and enhance human-machine interactions, deep learning (DL) models should learn how to master three fundamental questions: predict class labels for new inputs (the “What is the diagnosis for the X-ray shown in Figure 1?”), simulate changes in the situation to evaluate how this impacts class predictions (the “How would the absence of collapsed lung impact the presence of pneumothorax?”), and imagine alternative scenarios that would result in different class predictions (the “Why is the patient not having pneumothorax?”)—also known as counterfactual explanations (Wachter et al., 2017). Currently, no existing DL model is designed to answer all these questions by design.
Concept Bottleneck Models (CBMs) (Koh et al., 2020) are DL architectures designed to answer both “what?” and “how?” questions. They first predict human-understandable concepts (Kim et al., 2018), allowing experts to evaluate how intervening on such concepts influences the model’s class predictions. However, CBMs fall short in answering “why not?” queries as they are not trained to produce concept-based counterfactuals. More broadly, counterfactuals have been explored in concept-based models, e.g., for concept discovery (Stammer et al., 2024), while causal inference addresses this statistically (Louizos et al., 2017; Xia et al., 2022), seeking to discover and/or exploit the true causal relationships in the data. In contrast, we focus on addressing these questions from the model’s perspective, shedding light on its decision-making process. From this viewpoint, state-of-the-art counterfactual generation methods (Wachter et al., 2017; Pawelczyk et al., 2020; Guyomard et al., 2023; Nemirovsky et al., 2021) answer “what?” and “why not?” but not “how?”, especially when working with unstructured data such as images. Moreover, in these cases, counterfactual methods may generate an image that either i) looks nearly identical to the original, or ii) incorporates pixel-level modifications of difficult interpretation from the human perspective across thousands of pixels throughout the entire image (as shown in Figure 1). This scattering of modifications makes it hard to understand the answers to the “why not” questions. In addition, it is difficult to answer “how” questions because semantically coherent pixel-level interventions are hard to define and perform for domain experts (e.g., radiologists) (Abid et al., 2022). In summary, input-level counterfactuals often lack actionable insights, while concept-based models cannot answer counterfactual queries.
Contributions. To bridge these gaps, we introduce CounterFactual Concept Bottleneck Models (CF-CBM), a class of deep learning models designed to jointly master three fundamental questions for a classification problem - "What?", "How?" and "Why not?", generating interpretable concept-based counterfactuals. Our key innovation is a latent process that generates two similar concept vectors: one predicts the target class label (as in standard CBMs), while the other predicts an alternative class. Our experiments show that CF-CBMs achieve classification accuracy comparable to black-box models and existing CBMs (“What?”), rely on fewer important concepts leading to simpler explanations (“How?”), and produce interpretable, concept-based counterfactuals (“Why not?”). Additionally, training the counterfactual generator jointly with the CBM leads to two key benefits: (i) it alters the model’s decision-making, making it rely on fewer important concepts and (ii) it significantly increases the causal effect of concept interventions on class predictions, making the model more responsive to these changes. The code of this paper is publicly available [^1] https://github.com/gabriele-dominici/Counterfactual-CBM.
@@
Figure 2: Counterfactual CBM. For a given input sample, the task predictor answers “what?” queries predicting class labels. The concept predictor answers “how?” queries simulating changes in the scenarios through interventions. When taking the counterfactual latent distribution $z^{\prime}$, the concept predictor answers “why not?” queries via concept-based counterfactuals, which are easier to understand then counterfactuals at the input level.
@@ 2 Background
Concept Bottleneck Models (CBMs). Concept Bottleneck Models (Koh et al., 2020) are interpretable architectures that explain their predictions using high-level units of information (i.e., “concepts”). Given a sample’s raw features $x\in X\subseteq\mathbb{R}^{d}$ (e.g., an image’s pixels), a set of $r$ concepts $c_{i}\in C\subseteq\{0,1\}^{r}$ (e.g., “collapsed lung”, “visible pleural line”), and a set of $l$ class labels $y_{j}\in Y\subseteq\{0,1\}^{l}$ (e.g., labels “pneumothorax” or “no pneumothorax”). A CBM estimates the conditional distribution $\prod_{j}p(y_{j}\mid c_{1},\dots,c_{r})\prod_{i}p(c_{i}\mid x)$ where $p(c_{i}\mid x)$ is a set of independent Bernoulli distributions. At test time, human experts may intervene on mispredicted concept labels to improve CBMs’ task performance, simulating different scenarios. In a probabilistic perspective (Bahadori & Heckerman, 2020; Misino et al., 2022), (ideal) concepts $c$ represent key factors of variation (Kingma & Welling, 2013; Zarlenga et al., 2023) of observed data $x$ and $y$, as illustrated in the graphical model $x$$c$$y$
Counterfactual Explanations. A counterfactual is a hypothetical statement in contrast with actual events that helps us understand the potential consequences of different choices or circumstances  (Pearl et al., 2016). In the context of machine learning, Wachter et al. (2017) define counterfactual explanations as an optimization problem representing an answer to a “why not” question. The optimization aims to find for each observation $x$ the closest datapoint $x^{\prime}$ such that a classifier $m:X\rightarrow Y$ produces a class label $m(x^{\prime})$ that is different from the original label $m(x)$.
@@ 3 Counterfactual Concept Bottleneck Models
CounterFactual Concept Bottleneck Models (CF-CBMs, Figure 2) is a novel class of interpretable models designed to generate counterfactuals via variational inference at the concept level. In the following, we motivate CF-CBMs’ architecture and optimization objective (Section 3.1). Sections 3.3 and 3.2 show how CF-CBMs answer the different questions and how users can act upon generated counterfactuals.
@@ 3.1 Generating Concept-Based Counterfactuals
We also notice that multiple counterfactuals can be used to explain any given fact (Wachter et al., 2017; Pearl, 2000). To address this, an ideal solution would be to use a generative approach modeling the discrete distributions over $c$ and $c^{\prime}$. However, modeling such distributions directly might be unfeasible in practice (Richardson & Domingos, 2006) as it requires either modelling complex concept dependencies with discrete distributions which scale exponentially with the number of concepts, or assuming that concepts are independent of each other, an often unrealistic assumption. For this reason, we use a latent variable approach (Kingma & Welling, 2013) to model concept dependencies in a continuous latent distribution, which allows us to model discrete concept values as independent of each other given the latent variable.
Architecture. CF-CBMs are latent variable models generating counterfactuals via variational inference. To this end, they add two random variables $z$ and $z^{\prime}$ to Diagram 1. These variables represent latent factors of variation whose probability distributions are easier to model and sample compared to those for $c$ and $c^{\prime}$. We also include arrows from $c$ and $y$ to the counterfactual latent distribution $z^{\prime}$ in order to explicitly model the dependency of $z^{\prime}$ on the symbolic values of actual concept $c$ and class labels $y$, resulting in the following overall probabilistic graphical model:
@@
$p(c,y,z,c^{\prime},y^{\prime},z^{\prime},x)=p(c,y|z)p(c^{\prime},y^{\prime}|z^{\prime})p(x|z)p(\mathbf{z}|c,y)$ (3)
$p(c,y|z)=p(y|c)p(c|z),\quad p(c^{\prime},y^{\prime}|z^{\prime})=p(y^{\prime}|c^{\prime})p(c^{\prime}|z^{\prime}),\quad p(\mathbf{z}|c,y)=p(z)p(z^{\prime}|z,c,y)$ (4)
In our approach, $p(y|c)$ and $p(y^{\prime}|c^{\prime})$ are modeled as categorical distributions parametrized by the same task predictor $f$; $p(c|z)$ and $p(c^{\prime}|z^{\prime})$ as sets of Bernoulli distributions parametrized by the same concept predictor $\phi_{z}$. In practice, we assume that the input x is always observed at test time, making the p(x|z) term irrelevant. This allows $z$ and $z{\prime}$ to encode only the most salient information about the concepts c and $c{\prime}$, rather than all the details needed to reconstruct x. As a result, the optimization process is significantly simplified, enabling efficient counterfactual generation for complex data like images. Finally, $p(z)$ is a normal prior distribution and $p(z^{\prime}|z,c,y)$ is a learnable normal prior whose mean and variance are parametrized by neural networks $\phi_{p\mu}$ and $\phi_{p\sigma}$. The added complexity introduced with these additional black-box components is limited to the stages before concept extraction, enabling the encoding of information about possible counterfactuals. Thus, our model retains the interpretability of CBMs while providing the additional capability to explain decisions through counterfactuals.
@@
Amortized inference. CF-CBMs amortize inference needed for training by introducing two approximate Gaussian posteriors $q(z|x)$ and $q(z^{\prime}|z,c,y,y^{\prime})$ whose mean and variance are parametrized by a pair of neural networks $(\phi_{\mu},\phi_{\sigma})$ ($(\phi_{\mu^{\prime}},\phi_{\sigma^{\prime}})$, respectively). The corresponding inference graphical model (i.e. the encoder) is shown in Appendix A.
@@
Optimization problem. CF-CBMs are trained to optimize the log-likelihood of tuples $(c,y,y\textquoteright)$ ’ ), where $y\textquoteright$ ’ is randomly sampled from a uniform distribution for each observation of $x$, simulating a scenario in which a user requests the generation of a counterfactual for a specific label. Following a variational inference approach, we optimize the evidence lower bound of the log-likelihood, which results (see Appendix A) in the following objective function to maximize:
@@
$\mathcal{L}=$ $\overbrace{\lambda_{1}\mathbb{E}_{z\sim q(z|x)}[\log p(c|z)]+\log p(y|c)}^{\text{ reconstruction of $c$ and $y$}}-\overbrace{\lambda_{2}D_{KL}[q(z|x)||p(z)]}^{\text{ prior regularization on $z$}}$
@@
$+$ $\overbrace{\lambda_{3}\mathbb{E}_{z,z^{\prime},c^{\prime}\sim p(c^{\prime}|z^{\prime})q(z^{\prime}|\alpha)q(z|x))}[\log p(y^{\prime}|c^{\prime})]}^{\text{ reconstruction of $y^{\prime}$}}-\overbrace{\lambda_{4}D_{KL}[q(z^{\prime}|\alpha)||p(z^{\prime}|z,c,y)]}^{\text{prior regularization on $z^{\prime}$ }}$
@@
where $D_{KL}$ is the Kullback–Leibler divergence, $\alpha=(z,c,y,y^{\prime})$ and $\lambda_{i}$ are hyperparameters. Moreover, in order to enforce concept-based counterfactuals to be as close as possible to the current concept labels, we add an additional term to the objective (Kingma & Welling, 2022):
$\mathcal{L}_{dz}=-\overbrace{\lambda_{5}D_{KL}\left[q(z|x)||q(z^{\prime}|\alpha)\right]}^{\text{posterior distance}}-\overbrace{\lambda_{6}D_{KL}\left[p(z)||p(z^{\prime}|z,c,y)\right]}^{\text{prior distance}}$ (5)
3.2 Answering “what?”, “how?” and “why not?” queries
@@
CF-CBMs design allows them to answer “what?”, “how?” and “why not?” queries through the following steps (Figure 2):
1. What? Predict concept and class labels:
@@
    Sample from latent concept posterior: $z\sim\mathcal{N}(\phi_{\mu}(x),\phi_{\sigma}(x))$
@@
    Sample to predict concept and class labels: $\hat{y}\sim\text{Cat}(f(\hat{c})),\ \hat{c}\sim\text{Ber}(\phi_{z}(z))$
2. How? Simulate changes in the situation and evaluate the impact on the outcome. Similar to standard CBMs, it is possible to intervene at the concept level—since these concepts are inherently interpretable—and analyze how these modifications impact the model’s prediction:
  1. (a)
    Simulate changes intervening on concept values: $\hat{c}_{i}\mathrel{:}=\tilde{c}_{i}\quad\forall i\in\mathcal{I}$ where $\tilde{c}_{i}$ is the intervened value for each concept $i$ in the set $\mathcal{I}$ of the intervened concepts.
  2. (b)
    Sample to predict class labels based on the intervened concepts: $\tilde{y}\sim\text{Cat}(f(\hat{c}))$
3. Why not? Imagine interpretable counterfactuals via the generation of new concept tuples: